One canned juice drink is 20% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 15L that is 15% orange juice? How many liters of 20% and how many liters of 5%?
Answer by checkley77(12569)

.20x+.05(15-x)=.15*15
.20x+.75-.05x=2.25
.15x=2.25-.75
.15x=1.5
x=1.5/.15
x=10 liters of 20% orangw juice.
15-10=5 liters of 5% orange juice.
Proof:
.20*10+.05*5=.15*15
2+.25=2.25
2.25=2.25